# Diving into Self-Evolving Training for Multimodal Reasoning

### v1

@@ 4.1 Monitoring the Training Dynamics
Intuitively, two critical conditions must be met for the success of self-evolving training: (1) the presence of high-quality candidate responses generated by the model, otherwise self-evolving will not work no matter how strong the reward is; and (2) the reward function’s ability to effectively distinguish and prioritize these high-quality responses. These conditions align with the traditional reinforcement learning concepts of exploration and exploitation. Apparently, both exploration and exploitation capabilities are dynamic targets in self-evolving training, as the policy model evolves and the distribution of rollout responses changes with each iteration. To better understand these training dynamics, we propose tracking and visualizing four metrics:
@@
(Pass@K - Greedy) Accuracy: the difference between Pass@K and Greedy accuracy. Typically, Pass@K is an upper bound of Greedy Accuracy, and the gap roughly reflects the percentage of samples where the model, while failing in greedy decoding, can generate a correct response when sampling more candidates. This gap is crucial for the success of self-evolving training—a zero gap indicates that the model fails to explore correct responses for the current failure cases, suggesting that further training is unlikely to yield significant improvement.
Reward-Pass@2: the percentage of samples for which there exist correct responses among the top 2 responses ranked by the reward model. This metric directly reflects the exploitation efficacy of the reward model for the current policy. We choose Pass@2 since our training strategy involves selecting the top 2 responses using the reward model (§3.3).
@@ Results:
Figure 3 shows a clear trend where, as training progresses, the Pass@K metric continuously declines while greedy accuracy improves. This pattern indicates the loss of exploration ability, which hampers the model’s potential for continuous improvement and may lead to performance saturation. These observations are consistent with findings in text-only settings as reported by Wu et al. (2024). In Figure 4(a) we analyze Pass@K accuracy at various temperatures and observe a significant trend: despite a general decay in exploration ability, larger temperatures tend to resist this decline more effectively, allowing the model to maintain a stronger ability to explore in the mid to late stages of training. This observation suggests that the optimal temperature for training may need to be dynamically adjusted throughout the self-evolving process, rather than being fixed at the outset as is currently common practice. In Figure 4(b) we plot the (Pass@K - Greedy) accuracy, this phenomenon becomes even more pronounced, indicating that the model’s exploration during training is converging to greedy decoding. Additionally in Figure 4(c), we observe that the Reward-Pass@2 metric initially increases but quickly reaches a plateau, indicating that the reward model’s capacity to exploit further diminishes as training progresses. This limitation could be due to both the reduced exploration ability and the inherent constraints of the reward model. Next, we fix the reward model as a control variable and ask, how can we enhance exploration to allow the reward model to exploit more effectively?.222While improvements to the reward model could also enhance Reward-Pass@2, we reserve it for future work.
Reward-Pass@2 closely relates to the effectiveness of our self-evolving training strategy since our method selects top responses ranked by the reward model, and Reward-Pass@K directly reflects the quality of these 2 responses.333We note that there is a slight mismatch between Reward-Pass@2 and our training strategy, as we pre-filter responses using the ground-truth answer before the reward model reranks them. Ideally, a more aligned metric would measure the CoT reasoning quality of the top 2 responses, both containing correct answers. Given that there is no reliable method to score the quality of the thought processes, we consider Reward-Pass@2 as a reasonable approximation which turns out to be effective empirically. While Reward-Pass@2 naturally measures exploitation when the policy is fixed, the absolute value of this metric actually encapsulates both exploration and exploitation – its value would be low if the model fails to explore high-quality candidates.
Therefore, we hypothesize that enhancing the Reward-Pass@K scores for the current iteration through varied configurations could potentially improve the efficacy of self-evolving training. We fix reward model as a control variable and focus on modifying the model’s exploration capabilities to achieve this objective. Analysis in §4.1 suggests that the temperature, which is crucial for exploration, may require dynamic adjustment. Thus we propose to adjust the temperature automatically at each iteration based on the validation Reward-Pass@2 scores. This aims to optimize exploration so that the selected responses are of higher quality, potentially enhancing overall training effectiveness.
Full Results
Table 4 presents the results of our final approach as well as the comparison with representative baselines for all three models. We also demonstrate the performance on all sub-tasks of MathVista for a detailed and fine-grained analysis. We can see that by incorporating the dynamics of Reward-Pass@2, which balances both exploration and exploitation, our final recipe achieves the highest results for all of the 3 backbone LMMs and the improvement is generally consistent across all sub-tasks. In addition to overall trend, we observe that self-evolving training based on larger models yields more comprehensive improvements across various sub-tasks. We assume that the smaller model like InternVL2-2B may struggle to generalize its learned abilities across different domains as effectively as the larger models, such as MiniCPMV-2.5 and Phi-3.5-vision.
We also plot how the Pass@K and Reward-Pass@2 change for M-STaR (Reward-Pass@2) when trained on MiniCPM-V-2.5. To align with training, we show the metrics corresponding to the selected temperature in each iteration (see Appendix F for others). Figure 5 shows that compared with static strategy to choose a fixed temperature over the whole process, tuning it automatically mitigate the regression of Pass@K to help maintain the exploration ability. Besides, the Reward-Pass@2 is also generally higher than before. These further highlight the necessity to monitor the dynamics during training and adjust accordingly.
M-STAR on More Diverse Benchmarks
To further investigate how well M-STaR generalizes to benchmarks other than MathVista along, we select four extra multi-modal benchmarks focus on reasoning as well: M3CoT (Chen et al., 2024b), MMStar (Chen et al., 2024a), MMBench (Dev set, v1.1) (Liu et al., 2025), AI2D (Kembhavi et al., 2016). For MMStar and MMBench, we remove the perception sub-tasks in them to construct subsets focus more on reasoning. As shown in Table 5, models self-evolved with M-STaR consistently outperform both the base models and those trained with warmup across nearly all benchmarks. The only exception is InternVL2-2B, which underperforms on two benchmarks, aligning with the findings and speculations discussed above. Smaller models face greater challenges in generalizing beyond their training data, particularly on perception-intensive benchmarks like MMBench-R and AI2D. In contrast, larger models such as Phi-3.5-vision and MiniCPM-V-2.5 demonstrate significantly improved generalization, despite being trained with the same query set.
@@ 5 Conclusion
We dive into the self-evolving training for multimodal reasoning. Three static components are identified at first, namely the training method, reward model and the prompt variation. Through controlled experiments, we conclude a set of optimal design choices. On the other direction, we go deeper into the dynamics of self-evolving training to analyze the trade-off between exploitation and exploration. By monitoring the dynamics and adjusting key hyperparameters accordingly, we are able to further improve the model performance. We hope our work can provide insights and guidance for future research on self-evolving training for multimodal reasoning.
@@ Appendix A Details of Selected LMMs
is a powerful, openly released LMM. MiniCPM-V-2.5 leverages LLaMA-3-8B (Meta, 2024) for its language model and SigLIP (Zhai et al., 2023) as its vision encoder, resulting in strong multimodal capabilities. Its performance on a wide range of multimodal benchmarks significantly surpasses previous openly released LMMs such as LLaVA (Liu et al., 2023; 2024a) and Qwen-VL (Bai et al., 2023).
@@
Appendix B Collecting Warmup Training Data with Chain-of-Thought
Since our base model typically outputs the answer directly when responding to multimodal reasoning questions, during the warmup phase, we added additional instructions along with the input question, requiring the model to output the rationale. The instructions used in this process are as follows:
@@ Appendix E Measuring Response Relativity
Appendix F More Results for M-STaR
@@
Appendix G More Related Works
@@
On the other hand, several works have explored iterative approaches for self-improvement. Notably, STaR (Zelikman et al., 2022), ReSTEM (Singh et al., 2023), and V-STaR (Hosseini et al., 2024) retrain their models from the original checkpoint after each iteration, while ReST (Gulcehre et al., 2023) continuously fine-tunes the model starting from the previous iteration’s checkpoint. Reinforcement Learning (RL) techniques also fit into this iterative category, offering an online mechanism that tightly couples exploration and exploitation. RL methods, such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024), are frequently applied to unlabeled data, using an additional reward model to evaluate the quality of generated responses. GRPO, in particular, streamlines the process by removing the value model from PPO and instead leverages in-batch comparison to estimate advantages across different rollouts, providing a more stable alternative.
@@
Currently, the most common approach to improve multimodal reasoning capabilities continues to be supervised fine-tuning (SFT). For example, G-LLaVA (Gao et al., 2023) augments existing geometry question-answering datasets to fine-tune the LLaVA-1.5 model (Liu et al., 2024a). Math-LLaVA (Shi et al., 2024) selects and augments data from larger multimodal question-answer datasets, carefully balancing the difficulty of samples. Similarly, MAVIS (Zhang et al., 2024) focuses on the geometry and function domains and generates instruction-based tuning data through synthetic data engines.

### v3

@@ 4.1 Monitoring the Training Dynamics
Intuitively, two critical conditions must be met for the success of self-evolving training: (1) the presence of high-quality candidate responses generated by the model, otherwise self-evolving will not work no matter how strong the reward is; and (2) the reward function’s ability to effectively distinguish and prioritize these high-quality responses. These conditions align with the traditional reinforcement learning concepts of exploration and exploitation. Apparently, both exploration and exploitation capabilities are dynamic targets in self-evolving training, as the policy model evolves and the distribution of rollout responses changes with each iteration. To better understand these training dynamics, we propose tracking and visualizing three metrics, where we introduce a novel metric, Reward-Pass@2, to monitor the exploration-expolitation trade-offs:
@@
Reward-Pass@2: the ratio (%) of samples for which there exist correct responses among the top 2 responses ranked by the reward model. This metric directly reflects the exploitation efficacy of the reward model for the current policy. We choose Pass@2 since our training strategy involves selecting the top 2 responses using the reward model (§3.3).
@@
Results Figure 3 shows a clear trend where, as training progresses, the Pass@K metric continuously declines while greedy accuracy improves. This pattern indicates the loss of exploration ability, which hampers the model’s potential for continuous improvement and may lead to performance saturation. These observations are consistent with findings in text-only settings as reported by Wu et al. (2024). In Figure 4(a) we analyze Pass@K accuracy at various temperatures and observe a significant trend: despite a general decay in exploration ability, larger temperatures tend to resist this decline more effectively, allowing the model to maintain a stronger ability to explore in the mid to late stages of training. This observation suggests that the optimal temperature for training may need to be dynamically adjusted throughout the self-evolving process, rather than being fixed at the outset as is currently common practice.
| Variant | CPM | Phi | Intern |
| Base | 52.4 | 46.5 | 46.4 |
| +warmup | 52.8 | 49.3 | 47.6 |
| SFT | 54.7 | 49.5 | 41.9 |
| Iterative RFT | 55.7 | 50.2 | 47.5 |
| 55.1 | 50.5 | 47.9 | |
| Cont. Self-Evolving | 57.2 | 51.1 | 48.4 |
| + PRM Re-Rank | |||
| M-STaR (Reward-Pass@2) |
In Figure 4(b), we observe that the Reward-Pass@2 metric initially increases but quickly reaches a plateau, indicating that the reward model’s capacity to exploit further diminishes as training progresses. This limitation could be due to both the reduced exploration ability and the inherent constraints of the reward model. Next, we fix the reward model as a control variable and ask, how can we enhance exploration to allow the reward model to exploit more effectively?.111While improvements to the reward model could also enhance Reward-Pass@2, we reserve it for future work.
@@ 4.2 M-STaR– Final Recipe with Optimal Design Choices & Adaptive Explorations
Reward-Pass@2 closely relates to the effectiveness of our self-evolving training strategy since our method selects top responses ranked by the reward model, and Reward-Pass@K directly reflects the quality of these 2 responses.222We note that there is a slight mismatch between Reward-Pass@2 and our training strategy, as we pre-filter responses using the ground-truth answer before the reward model reranks them. Ideally, a more aligned metric would measure the CoT reasoning quality of the top 2 responses, both containing correct answers. Given that there is no reliable method to score the quality of the thought processes, we consider Reward-Pass@2 as a reasonable approximation which turns out to be effective empirically. While Reward-Pass@2 naturally measures exploitation when the policy is fixed, the absolute value of this metric actually encapsulates both exploration and exploitation – its value would be low if the model fails to explore high-quality candidates. Therefore, we hypothesize that enhancing the Reward-Pass@K scores for the current iteration through varied configurations could potentially improve the efficacy of self-evolving training. We fix reward model as a control variable and focus on modifying the model’s exploration capabilities to achieve this objective. Analysis in §4.1 suggests that the temperature, which is crucial for exploration, may require dynamic adjustment. Thus we propose to adjust the temperature automatically at each iteration based on the validation Reward-Pass@2 scores. This aims to optimize exploration so that the selected responses are of higher quality, potentially enhancing overall training effectiveness.
Full Results Table 4 presents the results of our final approach as well as the comparison with representative baselines. We also demonstrate the scores on all sub-tasks of MathVista in Table 6, Appendix H. We see that by incorporating the dynamics of Reward-Pass@2, which balances both exploration and exploitation, our final recipe achieves the highest results for all three backbone LMMs. In addition to overall trend, we observe that self-evolving training based on larger models yields more comprehensive improvements. We assume that the smaller model like InternVL2-2B may struggle to generalize its learned abilities across different domains as effectively as the larger models, such as MiniCPMV-2.5 and Phi-3.5-vision.
We also plot how the Pass@K and Reward-Pass@2 change for M-STaR when trained on MiniCPM-V-2.5. To align with training, we show the metrics corresponding to the selected temperature in each iteration (see Appendix G for others). Figure 5 shows that compared with choosing a fixed temperature over the whole process statically, tuning it automatically mitigate the regression of Pass@K to avoid the exploration loss. Besides, the Reward-Pass@2 is also generally higher than before. These further highlight the necessity to monitor the dynamics and adjust accordingly.
M-STAR on More Diverse Benchmarks To further investigate how well M-STaR generalizes to multiple benchmarks, we select four extra multi-modal ones focus on reasoning as well: M3CoT (Chen et al., 2024b), MMStar (Chen et al., 2024a), MMBench (Dev set, v1.1) (Liu et al., 2025), AI2D (Kembhavi et al., 2016). For MMStar and MMBench, we remove the perception sub-tasks to construct subsets focus more on reasoning. As shown in Table 5, models self-evolved with M-STaR consistently outperform both the base models and those trained with warmup across nearly all benchmarks. The only exception is InternVL2-2B, which underperforms on two benchmarks, aligning with the speculations discussed above. Smaller models face challenges in generalizing beyond their training data, particularly on perception-intensive benchmarks like MMBench-R and AI2D. In contrast, larger models such as Phi-3.5-vision and MiniCPM-V-2.5 show significantly improved generalization, despite being trained with the same query set.
@@ 5 Conclusion
We dive into the self-evolving training for multimodal reasoning. Three static components are identified at first, namely the training method, reward model and the prompt variation. Through controlled experiments, we conclude a set of optimal design choices. On the other direction, we also go deeper into the dynamics of self-evolving training to analyze the trade-off between exploitation and exploration. By balancing the training dynamics, we are able to further improve its performance. We hope our work can provide insights and guidance for future research on self-evolving training for multimodal reasoning.
Acknowledgments
This project is partially supported by NSFC Grant 62306177.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Details of Selected LMMs
is a powerful, openly released LMM. MiniCPM-V-2.5 leverages LLaMA-3-8B (Meta, 2024) for its language model and SigLIP (Zhai et al., 2023) as its vision encoder, resulting in strong multimodal capabilities. Its performance on a wide range of multimodal benchmarks significantly surpasses previous openly released LMMs such as LLaVA (Liu et al., 2023, 2024a) and Qwen-VL (Bai et al., 2023).
@@
Appendix B Warm-Up Phase to Unlock the Chain-of-Thought (CoT) Capability of LMMs
In our preliminary experiments, we found that open-source LMMs would directly output the answer given the query, while struggling to produce detailed chain-of-thought (CoT) reasoning processes. This may originate from the the scarcity of high quality rationales in most existing multimodal SFT training datasets (Masry et al., 2022; Shi et al., 2024), which limits the ability of open-source LMMs to generate detailed, step-by-step reasoning. Self-evolving training, however, requires responses with varying intermediate steps to allow models to learn effectively from on-policy data. To address this issue, we initiate a warm-up phase to collect some CoT data from the model itself as the first step before self-evolving training. Instead of prompting the model to answer questions directly, we prompt it to generate intermediate reasoning steps for a given triplet (question, image, and answer) using the following instruction:
For each triplet, we ask models to rollout 16 samples with . We then filter out results where the final answers do not match the ground truth and sample 100K from the generated dataset to create a warm-up CoT dataset with correct answers. Finally, we fine-tune our models on this dataset, treating it as a standard RFT process. Our iterative self-evolving training process will then start from this model checkpoint after the warm-up training.
@@ Appendix E Measuring Response Relativity
Appendix F Extra Comparison between RM-selected Solutions and Others
As a complement to our analysis in §3.3, we additionally plot the number of reasoning steps (split by ) for the top-2 solutions re-ranked by our process reward model, as well as for the other solutions. We observe that the top-2 solutions typically have fewer reasoning steps, indicating that the reward model can effectively identify solutions with fewer irrelevant steps and prioritize more straightforward ones.
Appendix G More Results for M-STaR
@@
Appendix H Full Results of MathVista
We present the complete evaluation results on MathVista for all three of our selected models. The scores for each subtask in MathVista are reported in Table 6. As shown, M-STaR achieves improvements across all subtasks compared to the pre-evolved models, particularly on geometric problems and math word problems. This demonstrates its enhanced comprehensive multimodal reasoning ability across multiple aspects.
Appendix I More Related Works
@@
On the other hand, several works have explored iterative approaches for self-improvement. Notably, STaR (Zelikman et al., 2022), ReSTEM (Singh et al., 2023), and V-STaR (Hosseini et al., 2024) retrain their models from the original checkpoint after each iteration, while RAFT (Dong et al., 2023), ReST (Gulcehre et al., 2023) and (Zhang et al., 2024a)continuously fine-tune models starting from the previous iteration’s checkpoint. Reinforcement Learning (RL) techniques also fit into this iterative category, offering an online mechanism that tightly couples exploration and exploitation. RL methods, such as PPO (Schulman et al., 2017) and GRPO (Shao et al., 2024), are frequently applied to unlabeled data, using an additional reward model to evaluate the quality of generated responses. GRPO, in particular, streamlines the process by removing the value model from PPO and instead leverages in-batch comparison to estimate advantages across different rollouts, providing a more stable alternative.
@@
Currently, the most common approach to improve multimodal reasoning capabilities continues to be supervised fine-tuning (SFT). For example, G-LLaVA (Gao et al., 2023) augments existing geometry question-answering datasets to fine-tune the LLaVA-1.5 model (Liu et al., 2024a). Math-LLaVA (Shi et al., 2024) selects and augments data from larger multimodal question-answer datasets, carefully balancing the difficulty of samples. Similarly, MAVIS (Zhang et al., 2024b) focuses on the geometry and function domains and generates instruction-based tuning data through synthetic data engines.
@@
Connecting to Recent Rule-based RL
Recent work in online RL with rule-based rewards has demonstrated strong performance on complex reasoning tasks like MATH (DeepSeek-AI et al., 2025; Zeng et al., 2025a), suggesting that using PRMs in GRPO training may constrain improvement. While these approaches leverage rule-based rewards to enhance the reasoning capabilities of (M)LLMs, our M-STaR framework is compatible with these findings for two key reasons. First, our reward strategy is also rule-based: we filter out responses with incorrect answers before applying PRM, effectively using PRM as a reranker to select trajectories with the highest-quality reasoning steps. Second, our training setup differs significantly from approaches like DeepSeek-R1. While R1 applies policy gradients at every step using PRM (as in GRPO), which can introduce reward hacking issues, our STaR-like (Sun et al., 2024; Zeng et al., 2025b) method avoids step-wise policy gradients. Although RL with rule-based rewards provides excellent performance, we believe exploring how to effectively train PRMs tailored to more general scenarios (Chae et al., 2025), and integrating PRMs into online RL training, remain valuable directions for achieving even higher performance ceilings.